OpenStudy (anonymous):

9x+y=72 X-6y=63 what is the solution of the system

OpenStudy (anonymous):

solved by the substitution method

OpenStudy (anonymous):

-6x+y=0 9x-12y=2 Multiply 1st by 12 to get: -72x + 12y = 0 9x - 12y = 2 Add to solve for "x": -63x = 2 x = -(-/63) Substitute into y = 6x to solve for "y": y = 6(-2/63) = -12/63 Ordered Pair: (-2/63,-12/63)

OpenStudy (anonymous):

hope this help

OpenStudy (anonymous):

what is de answer